# The Automated but Risky Game: Modeling Agent-to-Agent Negotiations and Transactions in Consumer Markets

### v3

@@ 3.2 Main Results
Agents’ Negotiation Capability Scales with Model Size. The scaling law of LLM suggests that model capabilities generally improve with increasing parameter sizes (kaplan2020scaling; hoffmann2022training; bi2024deepseek; zhang2024scaling). Do LLMs’ negotiation capabilities also exhibit a similar scaling pattern in our setting? We design two experiments using the Qwen2.5-Instruct family across six parameter scales (0.5B to 32B): (1) We conduct an in-family tournament where all six Qwen2.5-Instruct variants compete against each other as both buyers and sellers; (2) We benchmark against DeepSeek-R1 (guo2025deepseek), one of the strongest negotiation models, and each Qwen2.5-Instruct variant competes against DeepSeek-R1 as both buyer and seller. As shown in Figure 5, we observe a clear $PRR$ scaling pattern that models with more parameters are able to obtain more discounts as the buyer agent and higher profits as the seller agent.
4 Understanding Economic Risks for Real-World Users
@@
Autonomous AI agents could potentially bring huge economic value to the users in many settings. However, they may also introduce systematic risks when being deployed at scale  (feliu2001intelligent; jablonowska2018consumer; rohden2023recommendation; deng2025ai; hammond2025multi; chen2025position). In this section, we discuss the potential risks when both buyers and sellers delegate their negotiations and decision-making to AI agents and how models’ capability gaps and anomalies may translate into tangible economic losses for real users.
4.1 From Model Capability Gap to Economic Loss
In Sections §3.2, we discuss the capability gap of different models and also the asymmetric influence of buyer versus seller agent roles. Although such performance gaps may seem expected in experiments, deploying such agents in consumer settings could systematically disadvantage users who rely on less capable models. In particular, we view these interactions as imbalanced games, where one party deploys a significantly stronger agent than the other. Whether a strong buyer faces a weak seller or vice versa, the party with the weaker agent suffers a strategic disadvantage. Thus, one crucial question emerges: How does this strategic disadvantage translate into quantifiable economic loss?
Buyer Seller $PRR_{\text{Buyer}}$(%) Avg Payment(\$) $\boldsymbol{\Delta}$ from Baseline (%) Impact
Strong vs. Strong
DeepSeek-R1 DeepSeek-R1 6.11 1,423,090 Baseline
Weak-Buyer vs. Strong-Seller
GPT-3.5 DeepSeek-R1 2.99 1,452,699 +2.09% Buyer overpays by 2.09%
Qwen-7B DeepSeek-R1 3.14 1,454,633 +2.09% Buyer overpays by 2.09%
Qwen-14B DeepSeek-R1 5.08 1,438,834 +1.10% Buyer overpays by 1.10%
Strong-Buyer vs. Weak-Seller
DeepSeek-R1 GPT-3.5 17.35 1,221,980 –14.13% Seller earns 14.13% less
DeepSeek-R1 Qwen-7B 13.09 1,314,796 –7.62% Seller earns 7.62% less
DeepSeek-R1 Qwen-14B 11.27 1,325,570 –6.94% Seller earns 6.94% less
@@ 4.2 From Model Anomaly to Financial Risks
Fully automated, agent-based negotiation systems are prone to various anomalies stemming from unstable decision-making and imperfect instruction following of their base LLMs (lan2025exploring; zhang2025agent; cemri2025multi). While such failures may seem trivial or expected in research settings, they pose tangible risks to users in real-world settings. In this section, we analyze four model behavioral anomalies, pinpoint the conditions that trigger them, and outline how they can be translated into real financial loss for users. The detailed mathematical formula for the following anomaly measurement can be found in Appendix E.2.
Metric DeepSeek-R1 DeepSeek-V3 gpt-4.1 o4-mini o3 GPT-4o-mini GPT-3.5 Qwen2.5-7B Qwen2.5-14B
Out-of-Budget Rate 1.69 0.53 2.18 2.98 2.73 0.36 6.25 11.76 4.78
Out-of-Wholesale Rate 0.50 0.87 0.71 0.31 0.46 1.79 5.75 7.91 2.14
@@
Table 3: Overall Out-of-Budget ($OBR$) and Out-of-Wholesale Rates ($OWR$) across models. Bold values indicate the best performance, underlined values denote the second-best.
@@
Figure 6: Heatmaps of the $OWR$ (left) from the perspective of buyer agents, and the $OBR$ (right) from the perspective of seller agents, across different budget types.
Constraint Violation. Consider scenarios where a user authorizes an AI agent to negotiate on their behalf with a fixed budget $\beta$. If the agent accepts a deal above the budget, the agent may overdraw the account or exceed the user’s willingness to pay. Similarly, a seller agent agreeing to prices below the seller’s cost $p_{w}$ incurs guaranteed losses. In our evaluation, we quantify such anomaly using two metrics: Out-of-Budget Rate ($OBR$) and Out-of-Wholesale Rate ($OWR$). As shown in Figure 6, for $OBR$, we find that models with stronger negotiation capabilities, such as the DeepSeek series and Latest Generation GPT Series, including GPT-4.1, o4-mini, o3 and GPT-4o-mini, generally respect budget constraints and reject infeasible deals as buyers. However, models like GPT-3.5 and Qwen-7B frequently breach constraints, accepting deals above their budget in over 10% of all cases. This issue occurs across different budget settings but becomes even more serious when the user has a relatively low budget, posing serious risks for users in bad financial situations.
For buyer agents, all models correctly adhere to the budget limits in retail and high-budget scenarios, achieving 0% $OBR$. When designing the budget range, we deliberately set a relatively low budget (below the cost) to test whether agents can reject offers instead of completing transactions where buyers spend over their budget or sellers sell at a loss to accommodate buyers’ low budget. On the left, Figure 6 shows that most sellers exhibit higher $OWR$ under low-budget scenarios compared to other budget scenarios, with Qwen2.5-7b reaching almost 18.5%. It is worth noticing that even o4-mini–otherwise flawless across all other budget levels– occasionally capitulates under extreme price pressure, agreeing to below-cost deals in the low-budget scenario. Such a result suggests that while not following instructions has been considered a common but trivial issue in many scenarios, it can pose serious financial risks to both the buyer and seller in real consumer settings.
@@
Figure 7: Left: Overpayment Rate ($OPR$) from perspective of buyer agents across all budgets; Right: Two examples of dialogue that occurs overpayment due to high-budget diclosure.
@@
Excessive payment. Our experiments uncover a surprising anomaly: in some cases, buyer agents pay more than the listed retail price. We quantify this behavior with Overpayment Rate (OPR)–the proportion of successful deals in which the final transaction price is higher than the retail price despite the buyer’s budget affording a lower amount. As shown by Figure 7 (Left), overpayment often occurs under high-budget settings. Except for the DeepSeek family and Latest Generation GPT Series (GPT-4.1, o4-mini and o3), every model overpays to some degree when buyers have large $\beta$ values. To further investigate this issue, we qualitatively examined a sample of the negotiation history. As illustrated in Figure 7 (Right), overpayment often occurs after sellers ask buyers to reveal their budget early in the conversation. Despite our system prompt explicitly instructing buyers not to disclose their budget unless strictly necessary, many buyer agents reveal their budget easily. Sellers then try to anchor their offers to the disclosed number, even when it is higher than the listing price, and buyers accept the inflated deal without any objection.
Negotiation Deadlock. Imagine a user who uses an API-based buyer agent for negotiation, expecting it to operate efficiently within reasonable bounds. A lay user would typically assume the agent will either reach a deal or end the negotiation in proper situations. However, in our experiment, we observe that agents might continue bargaining even when the seller has clearly stated a firm bottom line, leading to unnecessarily long negotiations. This behavior wastes computational resources and undermines the practical utility of automation. Here we name this issue as “Negotiation Deadlock” and formally define a negotiation deadlock as any dialogue that reaches the maximum number of rounds $T_{\text{max}}$ without a final agreement or explicit rejection.
Figure 8: Left: Deadlock Rate ($DLR$) from perspective of buyer agents, presenting both overall performance and budget-stratified breakdowns; Right: Example of dialogue that occurs negotiation deadlock due to buyer refuse to reject the deal.
@@
We qualitatively examined a range of real negotiation histories and found that most of the negotiation deadlocks are behavioral, arising when agents become overly fixated on continuing the negotiation. For example, buyer agents often obsessively pursue price reductions even after sellers state their minimum acceptable price (Figure 8 (Right)). To quantitatively investigate this issue, we manually analyzed all the negotiation history and calculated the Deadlock Rate ($DLR$) for each model. We found that this issue is particularly prevalent among weaker buyer models operating under low-budget conditions, especially Qwen2.5-7B (see heatmap in Figure 8 (Left)). Due to the inherent capability gaps, these models may struggle to recognize when further negotiation is futile or when rejecting an offer would be more optimal, resulting in unnecessary turn-taking and resource waste.
@@
Figure 9: Average $PRR_{\text{Buyer}}$ of all models across different budget settings. Buyer agents are more likely to negotiate better deals in low-budget settings.
@@
Early Settlement. When analyzing buyer agents under different budget constraints, we observe a notable phenomenon that may cause the buyer to overpay: as the budget increases, particularly at or above the retail price, models tend to accept the seller’s proposed price as soon as it is below the budget rather than striving for better prices. In contrast, lower budgets (i.e., below retail price) appear to stimulate stronger bargaining behaviors, resulting in higher average price reduction rates $PRR_{\text{Buyer}}$. As shown in Figure 9, $PRR_{\text{Buyer}}$ exhibits a clear downward trend as the buyer budget increases, with a gap of nearly 9% between the highest and lowest price reduction rates. In practical deployments, buyer agents may derive their negotiation strategy from user-provided financial context, such as account balances or spending limits. If higher available funds systematically reduce the agent’s bargaining effort, users with generous budgets could end up consistently overpaying, not due to market necessity but because the agent passively accepts prices without seeking better deals.
@@ 5 Conclusion
With the growing deployment of AI agents in consumer settings, automated agent-to-agent negotiations will become widespread. We present an experimental framework to explore potential risks in such interactions. Our analysis shows that these negotiations inherently favor stronger agents, causing financial losses for users with weaker agents. Additionally, anomalies in LLM behavior can translate into real economic harm. Our study underscores significant risks associated with automating consumer negotiations through LLM agents.
Appendix A Appendix
Appendix B Related work
B.1 AI Negotiations
Negotiation plays a key role in modern society, and many researchers have studied negotiations from different perspectives  (rubinstein1982perfect; nash2024non; hua2024game; mensfelt2024autoformalizing). Due to the potential business value and the rising capabilities of deep neural networks, researchers and practitioners have explored methods to build automated negotiation models (zhou2016multiagent; lewis2017deal; he2018decoupling; bakker2019rlboa). More recently, large language models (LLMs) have shown strong capabilities in contextual understanding and strategic generation, leading to a growing interest in prompt-based LLM agents for complex negotiation tasks (abdelnabi2024cooperation; schneider2024negotiating; bianchi2024well; shea2024ace; yang2024makes; yang2025fraud).
B.2 AI Agent in Consumer Settings
A growing body of research examines AI agents in consumer-facing contexts, focusing on trust, decision delegation, and behavioral responses. Prior work has studied how agent intelligence and anthropomorphism shape consumer trust (song2024exploring; zhao2025consumers), and how task type affects willingness to delegate decisions (frank2021ai; fan2022exploring; yao2025your). Chatbots and similar agents have also been explored as service intermediaries that influence consumer experience and perceived agency (chong2021ai). While these studies offer important insights, they largely view agents as passive advisors or interfaces. Recent work begins to explore more active roles: ACE (shea2024ace) introduces a negotiation training environment for LLM agents, and FishBargain (kong2025fishbargain) develops a seller-side bargaining agent for online flea markets. However, few studies systematically analyze how consumer-side agents negotiate with business agents, or how agent capabilities shape negotiation outcomes in real scenarios.
Appendix C Discussion

### v4

@@ 3.3 Agents’ Negotiation Capability Scales with Model Size
The scaling law of LLM suggests that model capabilities generally improve with increasing parameter count (Kaplan et al., 2020; Hoffmann et al., 2022; Bi et al., 2024; Zhang et al., 2024). Do negotiation capabilities also exhibit a similar scaling pattern in our setting? We design two experiments using the Qwen2.5-Instruct family across six parameter scales (0.5B to 32B): (1) We conduct an in-family tournament where all six Qwen2.5-Instruct variants compete against each other as both buyers and sellers; (2) We benchmark against our strongest negotiation model DeepSeek-R1 (Guo et al., 2025), where each Qwen2.5-Instruct variant competes against DeepSeek-R1 as both buyer and seller. As shown in Figure 3, we observe a clear $PRR$ scaling pattern that models with more parameters are able to obtain more discounts as the buyer agent and higher profits as the seller agent.
Figure 3: Qwen models with more parameters obtain better deals as both sellers and buyers when they are negotiating with each other (Top) and DeepSeek-R1 (Bottom).
3.4 Understanding the Negotiation Gap via Model Specifications and Common Benchmarks.
To investigate variation sources in negotiation capacity, we select six representative models and collect four model characteristics [^2] Data from model providers’ sites or papers: https://openai.com/index/hello-gpt-4o/; https://arxiv.org/abs/2501.12948; https://qwenlm.github.io/blog/qwen2.5-llm/. GPT-4o-mini params estimated via https://arxiv.org/abs/2412.19260. : model size (in billions of parameters), general task performance (MMLU (Hendrycks et al., 2020)), mathematical ability (MATH (Hendrycks et al., 2021)), and scientific ability (GPQA (Rein et al., 2024)). We create a Negotiation Capacity Score ($NCS$) by combining three negotiation-relevant metrics, $PRR_{\text{Buyer}}$, reverse of Seller Price Reduction ($1-PRR_{\text{Seller}}$), and $RP$, through z-score normalization followed by averaging. We then compute Pearson correlations between each model’s $NCS$ and the four benchmark scores. As shown in Figure 4, negotiation capacity shows very strong correlation with general task performance on MMLU ($r=0.93$), along with substantial correlations with mathematical ($r=0.87$) and scientific ability ($r=0.80$). The weakest correlation appears with model size ($r=0.53$), likely because some high-parameter models belong to earlier generations with less optimized architectures, while exact parameter counts for commercial models are unavailable. This gap implies that in real-world scenarios where parties use models with different capabilities, one party would suffer economic losses (see case study in Appendix F.2).
4 Anomaly-Induced High-Stakes Risks
@@
Autonomous AI agents could potentially bring huge economic value to the users in many settings. However, they may also introduce systematic risks when being deployed at large  (Feliu, 2001; Jabłonowska et al., 2018; Rohden and Zeferino, 2023; Deng et al., 2025; Hammond et al., 2025; Chen et al., 2025). In this section, we discuss the potential risks when both buyers and sellers delegate their negotiations and decision-making to AI agents and how models’ anomalies may translate into tangible economic losses for real users. We also outline a potential method toward mitigating these risks.
@@ 4.1 From Model Anomaly to Financial Risks
Fully automated, agent-based negotiation systems are prone to various anomalies stemming from unstable decision-making and imperfect instruction following of their base LLMs (Lan et al., 2025; Zhang et al., 2025; Cemri et al., 2025). While such failures may seem trivial or expected in research settings, they pose tangible risks to users in real-world settings. In this section, we analyze four model behavioral anomalies, pinpoint the conditions that trigger them, and outline how they can be translated into real financial loss for users. The detailed mathematical formula for the following anomaly measurement can be found in Appendix D.2.
Model Out-of-Budget Out-of-Wholesale
DeepSeek-R1 (Guo et al., 2025) 1.69 0.50
DeepSeek-V3 (Liu et al., 2024) 0.53 0.87
gpt-4.1 (Hurst et al., 2024) 2.18 0.71
o4-mini (Hurst et al., 2024) 2.98 0.31
o3 (Hurst et al., 2024) 2.73 0.46
GPT-4o-mini (Hurst et al., 2024) 0.36 1.79
GPT-3.5 (Hurst et al., 2024) 6.25 5.75
Qwen2.5-7B (Yang et al., 2024a) 11.76 7.91
Qwen2.5-14B (Yang et al., 2024a) 4.78 2.14
@@
Figure 5: Overall Out-of-Budget (OBROBR) and Out-of-Wholesale Rates (OWROWR) across models. Bold = best, underline = second-best.
@@
Figure 6: Top: Heatmaps of the OWROWR from the perspective of buyer agents; Bottom: Heatmap of OBROBR from the perspective of seller agents, across different budget types.
@@
Figure 7: Left: Two examples of dialogue that occurs overpayment due to high-budget diclosure; Right: Example of dialogue that occurs negotiation deadlock due to buyer refuse to reject the deal.
Constraint Violation.
Consider scenarios where a user authorizes an AI agent to negotiate with a fixed budget $\beta$. If the agent accepts a deal above the budget, it may overdraw the account or exceed the user’s willingness to pay. Similarly, a seller agent agreeing to prices below cost $p_{w}$ incurs guaranteed losses. We quantify such anomalies using Out-of-Budget Rate ($OBR$) and Out-of-Wholesale Rate ($OWR$). As shown in Figure 6, models with stronger negotiation capabilities, such as DeepSeek series and Latest Generation GPT Series (GPT-4.1, o4-mini, o3, GPT-4o-mini), generally respect budget constraints and reject infeasible deals. However, models like GPT-3.5 and Qwen-7B frequently breach constraints, accepting deals above budget in over 10% of cases. This issue becomes more serious with low budgets, posing risks for users in poor financial situations.
For buyer agents, all models correctly adhere to budget limits in retail and high-budget scenarios, achieving 0% $OBR$. When designing the budget range, we deliberately set low budgets (below cost) to test whether agents can reject offers instead of completing transactions where buyers overspend or sellers sell at a loss. Figure 6 shows that most sellers exhibit higher $OWR$ under low-budget scenarios, with Qwen2.5-7b reaching almost 18.5%. Notably, even o4-mini, otherwise flawless across all other budget levels, occasionally capitulates under extreme price pressure, agreeing to below-cost deals in low-budget scenarios. This suggests that while instruction-following failures are often considered trivial, they can pose serious financial risks to both buyers and sellers in real consumer settings.
Excessive payment.
@@
Our experiments uncover a surprising anomaly: buyer agents sometimes pay more than the listed retail price. We quantify this behavior with Overpayment Rate (OPR), the proportion of successful deals where the final transaction price exceeds the retail price despite the buyer’s budget allowing a lower amount. As shown in Figure 8 (top), overpayment frequently occurs under high-budget settings. Except for the DeepSeek family and Latest Generation GPT Series (GPT-4.1, o4-mini and o3), every model overpays when buyers have large $\beta$ values. We qualitatively examine negotiation histories and found that overpayment often occurs after sellers ask buyers to reveal their budget early in conversations. Despite our system prompt explicitly instructing buyers not to disclose their budget unless necessary, many buyer agents reveal their budget easily. Sellers then anchor offers to the disclosed number, even when higher than the listing price, and buyers accept the inflated deal without objection.
Figure 8: Top: Overpayment Rate ($OPR$) from perspective of buyer agents across all budgets;Bottom: Deadlock Rate ($DLR$) from perspective of buyer agents across all budgets;
Negotiation Deadlock.
@@
Imagine a user who uses an API-based buyer agent for negotiation, expecting it to operate efficiently within reasonable bounds. Since users are billed per token or API call, they assume the agent will either reach a deal or end negotiation appropriately. However, we observe that agents might continue bargaining even when sellers state firm bottom lines, leading to unnecessarily long negotiations. This wastes computational resources and undermines automation’s practical utility. We define this issue as "Negotiation Deadlock," formally defined as any dialogue reaching maximum rounds $T_{\text{max}}$ without final agreement or explicit rejection. We qualitatively examined negotiation histories and found most deadlocks are behavioral, arising when agents become overly fixated on continuing negotiation. For example, buyer agents often obsessively pursue price reductions after sellers state minimum acceptable prices (Figure 8 (bottom)). To investigate quantitatively, we manually analyzed all negotiation histories and calculated Deadlock Rate ($DLR$) for each model. This issue is particularly prevalent among weaker buyer models under low-budget conditions, especially Qwen2.5-7B (see heatmap in Figure 7(right)). Due to capability gaps, these models struggle to recognize when further negotiation is futile or when rejecting offers would be optimal, resulting in unnecessary turn-taking and resource waste.
@@
Figure 9: Average $PRR_{\text{Buyer}}$ of all models across different budget settings.
Early Settlement.
@@
When analyzing buyer agents under different budget constraints, we observe a notable phenomenon that may cause overpayment: as budgets increase, particularly at or above retail price, models tend to accept sellers’ proposed prices once they fall below budget rather than striving for better deals. In contrast, lower budgets (below retail price) stimulate stronger bargaining behaviors, resulting in higher average price reduction rates $PRR_{\text{Buyer}}$. As shown in Figure 10, $PRR_{\text{Buyer}}$ exhibits a clear downward trend as buyer budgets increase, with nearly 9% gap between highest and lowest price reduction rates. In practical deployments, buyer agents may derive negotiation strategies from user-provided financial context, such as account balances or spending limits. If higher available funds systematically reduce agents’ bargaining effort, users with generous budgets could consistently overpay, not due to market necessity but because agents passively accept prices without seeking better deals.
4.2 Anomaly Mitigation via RL-based Prompt Optimization
Figure 10: Online Mulit-armed Bandit Prompt Optimization for anomalies mitigation
To mitigate negotiation anomalies, we experiment with Qwen2.5-7B, the buyer model with the highest anomaly rate. The goal is to find prompts that reduce overpayment, out-of-budget transactions, and deadlocks. We formulate prompt search as an online multi-armed bandit RL problem over 96 candidate prompts generated by combining strategy options (e.g., budget emphasis, price-increase policy, progress threshold). The policy over $M$ arms is softmax: $\pi_{i}(\theta)=\frac{e^{\theta_{i}}}{\sum_{j}e^{\theta_{j}}}$, actions $a_{t}\sim\pi(\cdot|\theta)$, reward $r_{t}$ compared to baseline $b_{t}$, and only the chosen arm is updated: $\theta_{a_{t}}\leftarrow\theta_{a_{t}}+\eta(r_{t}-b_{t})(1-\pi_{a_{t}})$. Rewards penalize high-budget overpayment, low-budget out-of-budget (OOB), and deadlocks. The reward penalizes undesirable behaviors: high-budget overpayment, low-budget out-of-budget transactions, and negotiation deadlocks. Formally,
$r_{t}=-w_{1}\cdot\mathbbm{1}[b_{high}\land overpay]\\ -w_{2}\cdot\mathbbm{1}[b_{low}\land oob]-w_{3}\cdot\mathbbm{1}[deadlock],$ (1)
where $w_{i}>0$ are penalty weights. The optimal prompt is selected as the arm with the largest $\theta_{i}$ after training. Detailed training setting can refers to Appendix G. Overall, from Table 2, prompt optimization effectively reduces out of budget errors, while overpayment and deadlock are harder to mitigate. This preliminary result demonstrates the potential of RL-based prompt tuning to improve negotiation safety and inspires future research in secure AI agent deployment.
Anomaly Vanilla Online Bandit
Out of Budget (↓) 18.4 1.3
Overpay (↓) 8.1 8.3
Deadlock (↓) 4.0 4.0
Table 2: Effect of online bandit-based prompt optimization on negotiation anomalies (%).
@@ 5 Conclusion
With large-scale deployment of AI agents in real consumer settings, agent-to-agent interactions will become ubiquitous in the near future. But what happens when we fully automate negotiation and deal-making with consumer and seller authorized AI agents? In this paper, we design an experimental framework to investigate potential issues and risks in agent-to-agent negotiations and transactions. Our analysis reveals that agent-to-agent negotiation is naturally an imbalanced game where users with less capable agents face significant financial loss against stronger agents. Furthermore, we found that LLMs’ anomalies might transfer to real economic loss when deployed in real consumer settings. Our paper highlights the potential risks of using LLM agents to automate negotiation and transactions in real consumer settings.
6 Limitation
While this work primarily focuses on evaluating risks and performance disparities in fully delegated agent-to-agent negotiation and transaction scenarios, it does not provide a complete account of systematic mitigation strategies. Our mitigation experiment, an RL-based prompt optimization method, demonstrates the potential of reinforcement learning for reducing anomalies, but remain preliminary in scope. Future research should thus go beyond diagnosis toward jointly optimizing negotiation performance and risk reduction, ideally within real-world, human-in-the-loop platforms.
Appendix A Related work
A.1 AI Negotiations
Early research on negotiation was rooted in game theory, with foundational frameworks such as the alternating offers model (Rubinstein, 1982) and Nash’s non-cooperative game theory (Nash, 2024) forming the basis for subsequent AI negotiation studies (Hua et al., 2024; Mensfelt et al., 2024). With advances in deep learning, researchers developed negotiation models using supervised and reinforcement learning (Zhou et al., 2016; Lewis et al., 2017; He et al., 2018; Bakker et al., 2019). More recently, large language models (LLMs) have shown strong capabilities in contextual understanding and strategic generation, leading to a growing interest in prompt-based LLM agents for complex negotiation tasks (Abdelnabi et al., 2024; Schneider et al., 2024; Bianchi et al., 2024; Shea et al., 2024; Yang et al., 2024b, 2025).
A.2 AI Agent in Consumer Settings
A growing body of research examines AI agents in consumer-facing contexts, focusing on trust, decision delegation, and behavioral responses. Prior work has studied how agent intelligence and anthropomorphism shape consumer trust (Song and Lin, 2024; Zhao et al., 2025), and how task type affects willingness to delegate decisions (Frank et al., 2021; Fan and Liu, 2022; Yao et al., 2025). Chatbots and similar agents have also been explored as service intermediaries that influence consumer experience and perceived agency (Chong et al., 2021). While these studies offer important insights, they largely view agents as passive advisors or interfaces. Recent work begins to explore more active roles: ACE (Shea et al., 2024) introduces a negotiation training environment for LLM agents, and FishBargain (Kong et al., 2025) develops a seller-side bargaining agent for online flea markets. However, few research systematically analyzes how consumer-side agents negotiate with business agents, or how agent capabilities shape negotiation outcomes in real scenarios. Our work aims to address this gap.
Appendix B Discussion